The ratio of girls to boys in a movie school is 5:6, on a rainy day 1/6 of girls and 1/4 of boys were absent. If total number of pupils present that day was 390. How many of the pupils were boys.


1
Expert's answer
2021-01-26T03:31:44-0500

Let number of girls in the school are "5x" .Then number of boys are "6x."

On the rainy day "\frac {1}{6}" of girls and "\frac{1}{4}" of boys were absent. Therefore "\frac{5}{6}" of girls and "\frac{3}{4}" of boys are present.

So number of girls present on that day are "=5x×(\frac{5}{6})=\frac{25x}{6}"

and number of boys present on that day are "=6x×(\frac{3}{4})=\frac{9x}{2}"

Therefore according to the problem, "\frac{25x}{6}+\frac{9x}{2}=390"

"\implies \frac{52x}{6}=390"

"\implies x=45"

"\therefore" Number of pupils were boys are "=6x=6\u00d745=270" .